OpenStudy (lisa123):

suppose x and y are integers with 2x+y=30. Find the minimum value of x^2 +y

OpenStudy (lisa123):

the answer is 29. I just dont understand why

OpenStudy (amistre64):

Ever hear of Lagrange multipliers?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

simple substitution and then find vertex

OpenStudy (lisa123):

nope

OpenStudy (zarkon):

\[y=30-2x\]\[ x^2+y =x^2-2x+30\] vertex at 1 \[1^2-2\cdot 1+30=29\]

OpenStudy (lisa123):

how did you get x^2+y =x^2-2x+30

OpenStudy (zarkon):

replace y with \[y=30-2x=-2x+30\]

OpenStudy (lisa123):

ohhh wow okay I understand!

OpenStudy (lisa123):

Thank you both @Zarkon @IrishBoy123!!!